The following data were collected from the gravime
Answer #1

Mass of hydrated salt can be found as

Mass of fired crucible, lid and hydrated salt - Mass of fired crucible & lid

Mass of hydrated salt for all the trials is calculated below

Trial 1 ------ 21.626 g - 19.437 g = 2.189 g

Trial 2 ------ 25.111 g - 20.687 g = 4.424 g

Trial 3 ------- 22.167 g - 18.431 g = 3.736 g

_________________________________________________________________________________________

Mass of anhydrous salt is calculated as

Mass of fired crucible, lid and anhydrous salt - Mass of fired crucible & lid

Mass of anhydrous salt for all the trials is calculated below

Trial 1 ----------- 21.441 g - 19.437 g = 2.004 g

Trial 2 ---------- 24.702 g - 20.687 g = 4.015 g

Trial 3 --------- 21.762 g - 18.431 g = 3.331 g

_________________________________________________________________________________________________

Mass of water lost is calculated as

Mass of hydrated compound - Mass of anhydrous compound

Mass of water lost for all the trials is calculated below

Trial 1 ------- 2.189 g - 2.004 g = 0.185 g

Trial 2 -------   4.424 g - 4.015 g = 0.409 g

Trial 3 -------   3.736 g - 3.331 g = 0.405 g

________________________________________________________________________________________________

Percent by mass of volatile water in hydrated compound is calculated as

Mass of water lost / Mass of hydrated compound * 100

Percent by mass of volatile water for all the trials is given below

Trial 1 ---------- 0.185 g / 2.189 g * 100 = 8.45%

Trial 2 -------- 0.409 g / 4.424 g * 100 = 9.26%

Trial 3 ---------- 0.405g / 3.736 g * 100 = 10.84 %

___________________________________________________________________________________________________

Average % H2O = 8.45 + 9.26 + 10.84 / 3 = 9.52 %

Average percent H2O in hydrated salt is 9.52%

__________________________________________________________________________________________________

Standard deviation is calculated as follows

Let us make a table to find (x- average x ) values

Trial x ( % of water) (x - average x) ( x - average x)^2
1 8.45 8.45 - 9.52 = -1.07 (-1.07)^2 = 1.1449
2 9.26 9.26 - 9.52 = -0.26 (-0.26)^2 = 0.0676
3 10.84 10.24 - 9.52 = 0.72 (0.72)^2 = 0.5184

Total ( x - average x)^2 = 1.1449 + 0.0676 + 0.5184 = 1.7309

n = number of trials, = 3

Substituting above values in S.D formula, we get

S.D . = Sq rt ( 1.7309 / 3-1)

S.D = Sq rt ( 0.8655)

S.D = 0.9303

Standard deviation of % of H2O is 0.9303

__________________________________________________________________________________________________

Relative standard deviation is calculated as

% RSD = Standard deviation / Mean( average of % H2O) * 100

% RSD = 0.9303 / 9.52 * 100

% RSD = 9.772 %